### v1

@@ 1 Introduction
In this work we explore the implementation of a tomographic, angular ADF+ARF analysis of a pseudo-spectroscopic survey like J-PLUS (Cenarro et al., 2019). By pseudo-spectroscopic surveys we refer to multi narrow/medium-width optical band surveys like COMBO (Wolf et al., 2003), ALHAMBRA (Moles et al., 2008), SHARDS (Pérez-González et al., 2013), PAU (Martí et al., 2014), J-PLUS (Cenarro et al., 2019), or J-PAS (Benitez et al., 2014), for which, in every pixel in the footprint, there exists a pseudo-spectrum with as many data points as the number of filters carried in the optical system. Typically, for these surveys the spectral resolution roughly equals the number of optical filters, $R\sim N_{\rm filter}$, with $N_{\rm filter}=17,20,12$, and $56$ for COMBO, ALHAMBRA, J-PLUS, and J-PAS, respectively. Such modest level of effective spectral resolution can not be compared to typical spectroscopic surveys (for which $R\gtrsim 500$), although it still leaves room for a wealth of science cases provided the photometric redshift precision can easily reach (and surpass) the 1 % level for a significant fraction of the detected galaxies. For instance, J-PAS with $N_{\rm filter}=56$ should still be sensitive to the radial BAO (Chaves-Montero et al., 2018), or should provide a superb catalogue of galaxy clusters and groups (Ascaso et al., 2016; Doubrawa et al., 2024; Maturi et al., 2023).
@@
In Sect. 2 we introduce and describe the J-PLUS survey, together with its third data release and the selection of the galaxy sample to analyze. In Sect. 3 we outline our approach to correct for observational systematics, and its outcome. In Sect. 4 we describe out theoretical model for the ADF and ARF angular power spectrum, and the parameters we attempt to constrain from their comparison to the actual power spectra measured in J-PLUS DR3 data. In Sect. 5 we present our results for the ADF-only, ARF-only, and the combined ADF$+$ARF analyses, with some further cross-correlation study of J-PLUS data with Planck’s CMB convergence $\kappa$ map. We discuss our findings in Sect. 6, and conclude in Sect. 7.
@@ 2 J-PLUS DR3
where $z_{\rm ml}$ is the redshift maximizing the posterior photo-$z$ PDF ${\cal P}(z|G)$ for a given source being a “galaxy” ($G$). High values of the odds parameter should hence correspond to “packed” or “self-contained” photo-$z$ PDFs around $z_{\rm best}$, or precise determination of the photo-$z$s.
@@ 2.2 The high odds galaxy sub-sample
The goal of this work is to conduct tomography of the J-PLUS DR3 galaxy angular density and redshift fields. This tomography will be conducted under Gaussian redshift shells of widths $\sigma_{z}=[0.01,0.03,0.05]$ and nominally centered in a regular redshift array starting at $z_{\rm min}=0.05$ and monotonically increasing with steps of $\Delta z=0.02$ up to $z_{\rm max}=0.39$. While this amounts to 18 different redshifts, we shall see that due to the impact of systematics the actual redshift range under analysis will restrict to $z\in[0.05,0.25]$ or 11 distinct central redshifts, which give rise to $11\times 3=33$ different Gaussian shells after accounting for the three widths under consideration.
@@ 5 Results
The joint ADF$+$ARF fits (red curves) are in many cases indistinguishable from the ADF- and ARF-only fits, particularly at low redshifts. Some tension between the 1- and 2-probe fits starts to appear at $z\gtrsim 0.17$, at a higher level for ARF than for ADF. Nevertheless, similar best-fit curves for one- and two-probes do not necessarily involve similar best-fit parameters. Indeed, as can be seen in Fig. 9 for the case where $\sigma_{\rm Err}$ is constant for all redshifts and the BANNJOS selection, the ADF-only and ARF-only contours (blue and green colours, respectively) are clearly incompatible for $\sigma_{\rm Err}$ (the first column of this plot corresponds to this parameter). The tension between the bias contours grows with redshift, surpassing the $\sim 2-\sigma$ level by first time at $z=0.17$, and then again at $z=0.23$ and $z=0.25$. The bias values from ADF at $z<0.17$ are systematically higher than those from ARF, but at the same time, $\sigma_{\rm Err}$ is also much higher. These two parameters are obviously correlated, since higher values of the bias $b_{g}(z)$ combined with higher values of $\sigma_{\rm Err}$ can yield similar amplitudes of the $C_{\ell}$s than those obtained under lower values of these two parameters, provided that each of them have opposite effect on the $C_{\ell}$ amplitudes. Again for the BANNJOS catalogue, Fig 10 shows the projected best-fit values for the parameters under the first parameter configuration (constant $\sigma_{\rm Err}$. The joint ADF$+$ARF fit (in green colour) is driven by the ARF on $\sigma_{\rm Err}$ (since ARF are significantly more sensitive to this parameter), and also for bias parameters below $z<0.19$. Above this redshift the ADF$+$ARF values are dominated by the ADF. The value provided for $\sigma_{\rm Err}$ by the photo-$z$ code LePhare is $\sigma_{\rm Err}=0.010$, which is below (at the $\simeq 9~{}\sigma$ level) the distribution for this parameter from ARF, and also clearly discrepant with the value provided by ADF. Regarding the velocity amplitude ${\cal A}_{v}$, there is a significant mismatch between the ADF and the ARF outputs: ARF see no evidence for radial peculiar velocities (${\cal A}_{v}=0.42\pm 0.57$), ADF point to ${\cal A}_{v}=3.4\pm 1.0$, while the joint ADF$+$ARF fit provides an intermediate solution that just by coincidence peaks around ${\cal A}_{v}\sim 1$, ${\cal A}_{v}=0.90\pm 0.56$.
@@
If we instead look at the results from the second configuration ($\sigma_{\rm Err}$ as a function of redshift), we find again for the BANNJOS catalogue that there is actually agreement for $\sigma_{\rm Err}$ at low redshifts between both probes, although systematically higher for ADF than for ARF, which again translates into higher ADF bias estimates. This is shown in the corner plot Fig. 11, whose results are summarised in Fig. 12. A careful look at Fig. 11 reveals a positive correlation of the $\{\sigma_{\rm Err}(z),b_{g}(z)\}$ parameters for the same shells: this is found in a diagonal whose first panel is located in the first column and 13th row, displaying the correlation between $\sigma{\rm Err}(z)$ and $b(z)$ for the first redshift bin. This diagonal follows further in the 14th row and second column. In Fig. 12 one can also notice that it is only at the highest redshift shells ($z>0.15$) when $\sigma_{\rm Err}$ estimates for ADF reach the $\sim 0.025$ level, which is higher than any ARF $\sigma_{\rm Err}$ estimate. Thus, on what regards $\sigma_{\rm Err}$ and $b_{g}(z)$, ADF and ARF agree for $z<0.17$.
@@ 7 Conclusions
The estimates of $\sigma_{\rm Err}(z)$ are compatible from either the ADF or the ARF probe for $z<0.17$ (but almost always higher for ADF than for ARF), a redshift range where we show systematics to have a low impact on the angular distribution of galaxies. These estimates of $\sigma_{\rm Err}(z)$ scatter around $\sigma_{\rm Err}(z)=0.014$ with a typical uncertainty of about 30 %, while the estimate from LePhare is $\langle\sigma_{\rm Err}(z)\rangle=0.010$. On that same redshift range, the galaxy bias estimates from both ADF and ARF seem to grow monotonically with $z$, with a very similar slope, from $b(z)\simeq 0.7$ at $z=0.05$ up to $b(z)\simeq 1.6$ at $z=0.15$. Although compatible, ADF bias estimates again lie systematically above ARF ones, and this is understood given the correlation between $b_{g}(z)$ and $\sigma_{\rm Err}(z)$ for any given redshift shell: higher values of these two parameters impact the amplitude of the angular power spectrum in opposite ways, and it is thus possible to leave the amplitude of the $C_{\ell}$s practically untouched after increasing both parameters conveniently.
@@ acknowledgements
C.H.-M. acknowledges the hospitality of the Centro de Estudios de Física del Cosmos de Aragón (ceFca), where part of this work took place. C.H.-M. also acknowledges the support of the Spanish Ministry of Science and Innovation project PID2021-126616NB-I00, and the contribution from the IAC High-Performance Computing support team and hardware facilities. VM thanks CNPq (Brazil) and FAPES (Brazil) for partial financial support. ET acknowledges funding from the HTM (grant TK202), ETAg (grant PRG1006) and the EU Horizon Europe (EXCOSM, grant No. 101159513). AE acknowledges the financial support from the Spanish Ministry of Science and Innovation and the European Union - NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA. Based on observations made with the JAST80 telescope at the Observatorio Astrofísico de Javalambre (OAJ), in Teruel, owned, managed, and operated by the Centro de Estudios de Física del Cosmos de Aragón. We acknowledge the OAJ Data Processing and Archiving Unit Department(DPAD, Cristóbal-Hornillos et al., 2012) for reducing and calibrating the OAJ data used in this work. Funding for the J-PLUS Project has been provided by the Governments of Spain and Aragón through the Fondo de Inversiones de Teruel; the Aragón Government through the Research Groups E96, E103, E16_17R, E16_20R, and E16_23R; the Spanish Ministry of Science, Innovation and Universities (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grants PID2021-124918NB-C41, PID2021-124918NB-C42, PID2021-124918NA-C43, PID2021-124918NB-C44, PGC2018-097585-B-C21 and PGC2018-097585-B-C22; the Spanish Ministry of Economy and Competitiveness (MINECO) under AYA2015-66211-C2-1-P, AYA2015-66211-C2-2, AYA2012-30789, and ICTS-2009-14; and European FEDER funding (FCDD10-4E-867, FCDD13-4E-2685). The Brazilian agencies FINEP, FAPESP, and the National Observatory of Brazil have also contributed to this project. Some of the results in this paper have been derived using the HEALPix (Górski et al., 2005) package.

### v2

@@ 1 Introduction
In this work we explore the implementation of a tomographic, angular ADF+ARF analysis of a pseudo-spectroscopic survey like J-PLUS (Cenarro et al., 2019). By pseudo-spectroscopic surveys we refer to multi narrow/medium-width optical band surveys like COMBO (Wolf et al., 2003), ALHAMBRA (Moles et al., 2008), SHARDS (Pérez-González et al., 2013), PAU (Martí et al., 2014b), J-PLUS (Cenarro et al., 2019), or J-PAS (Benitez et al., 2014), for which, in every pixel in the footprint, there exists a pseudo-spectrum with as many data points as the number of filters carried in the optical system. Typically, for these surveys the spectral resolution roughly equals the number of optical filters, $R\sim N_{\rm filter}$, with $N_{\rm filter}=17,20,12$, and $56$ for COMBO, ALHAMBRA, J-PLUS, and J-PAS, respectively. Such modest level of effective spectral resolution can not be compared to typical spectroscopic surveys (for which $R\gtrsim 500$), although it still leaves room for a wealth of science cases provided the photometric redshift precision can easily reach (and surpass) the 1 % level for a significant fraction of the detected galaxies. For instance, J-PAS with $N_{\rm filter}=56$ should still be sensitive to the radial BAO (Chaves-Montero et al., 2018), or should provide a superb catalogue of galaxy clusters and groups (Ascaso et al., 2016; Doubrawa et al., 2024; Maturi et al., 2023).
@@
In Sect. 2 we introduce and describe the J-PLUS survey, together with its third data release, its photo-$z$ estimation, and the selection of the galaxy sample to analyze. In Sect. 3 we outline our approach to correct for observational systematics, and its outcome. In Sect. 4 we describe out theoretical model for the ADF and ARF angular power spectrum, and the parameters we attempt to constrain from their comparison to the actual power spectra measured in J-PLUS DR3 data. In Sect. 5 we present our results for the ADF-only, ARF-only, and the combined ADF$+$ARF analyses, with some further cross-correlation study of J-PLUS data with Planck’s CMB convergence $\kappa$ map. We discuss our findings in Sect. 6, and conclude in Sect. 7.
@@ 2 J-PLUS DR3
where $z_{\rm ml}$ is the redshift maximizing the posterior photo-$z$ PDF ${\cal P}(z|G)$ for a given source being a “galaxy” ($G$). High values of the odds parameter should hence correspond to “packed” or “self-contained” photo-$z$ PDFs around $z_{\rm best}$, or precise determination of the photo-$z$s. Details on the photo-$z$ estimation follow next.
2.1 Photometric redshifts in J-PLUS DR3
The photo-$z$ of J-PLUS DR3 were obtained using jphotoz, a python package that is part of jype, the data reduction pipeline for J-PLUS and J-PAS (Vázquez Ramió et al, 2025). This pipeline acts as an interface between the J-PLUS database and the actual photo-$z$ engine(s) (which in the case of J-PLUS DR3 is LePhare, Arnouts & Ilbert, 2011). jphotoz handles all the pre-processing of the data that is fed to LePhare and the post-processing of its output for ingestion into the database.
The procedure for obtaining photo-$z$ for J-PLUS DR3 with jphotoz/LePhare is similar to that used for the miniJPAS survey, which is described in detail in Hernán-Caballero et al. (2021). Here we summarize the main steps and highlight the differences.
The photometry used is the PSF-corrected photometry (PSFCOR) from the dual-mode catalogs. PSFCOR magnitudes are corrected for Galactic extinction and re-calibrated tile by tile using stellar population synthesis models (see Sects. 2 and 3 in Hernán-Caballero et al., 2021). LePhare works by scanning a redshift range from $z_{\rm min}$ to $z_{\rm max}$ in steps of $z_{\rm step}$. For J-PLUS, $z_{\rm min}$ = 0.0, $z_{\rm max}$ = 1.0, and $z_{\rm step}$ = 0.005 were used. A likelihood distribution $\mathcal{L}$($z$) is obtained from the $\chi^{2}$ of the best-fitting template at each $z$. The template library, CEFCA_minijpas, contains 50 synthetic galaxy spectra generated with CIGALE (Boquien et al., 2019). A redshift prior $P(z,T|m)$, where $T$ denotes a galaxy spectra template and $m$ refers to $r$-band magnitude, is derived from galaxy counts in the VIMOS VLT Deep Survey (VVDS; Le Fèvre et al., 2005). Such prior is consequently applied to $\mathcal{L}$($z$) to obtain the redshift probability distribution $P(z)$. For further details on the configuration of LePhare, the templates, and the redshift prior, see Sect. 4 in Hernán-Caballero et al. (2021).
While the $P(z)$ thus obtained is meaningful only for galaxies, it is computed for all J-PLUS sources brighter than $r=22$ regardless of their classification. This allows to compensate for the classification uncertainty in statistical studies. The mode of $P(z)$ is taken as the best point estimate for the redshift of the source, $z_{\rm best}$. The odds parameter is defined as the integral of $P(z)$ in the range $z_{\rm best}\pm 0.03(1+z_{\rm best})$ and corresponds to the predicted probability of $|\Delta z|\equiv(|z_{\rm best}-z_{\rm true}|)/(1+z_{\rm true})$ being $<3~{}\%$.
In miniJPAS, a contrast correction was applied to the raw $P(z)$ to ensure that the fraction of galaxies with $|\Delta z|<3~{}\%$ matches their average $odds$ value in a spectroscopic sample where $z_{\rm true}$ is known. This correction was not applied for J-PLUS DR3 due to the lack of a representative spectroscopic sample. As a consequence, $odds$ values for J-PLUS DR3 can be overconfident.
@@ 2.3 The high odds galaxy sub-sample
The goal of this work is to conduct tomography of the J-PLUS DR3 galaxy angular density and redshift fields. This tomography will be conducted under Gaussian redshift shells of widths $\sigma_{z}=[0.01,0.03,0.05]$ and nominally centred on a regular redshift array starting at $z_{\rm min}=0.05$ and monotonically increasing with steps of $\Delta z=0.02$ up to $z_{\rm max}=0.39$. While this amounts to 18 different redshifts, we shall see that due to the impact of systematics the actual redshift range under analysis will restrict to $z\in[0.05,0.25]$ or 11 distinct central redshifts, which give rise to $11\times 3=33$ different Gaussian shells after accounting for the three widths under consideration. We remark that we choose a Gaussian shape for the redshift shells for sake of simplicity, since a priori we could have chosen any other shape. We also emphasize that galaxies belonging to a given redshift shell are selected according to their observed redshift, which includes errors associated in the photo-$z$ estimation.
@@ 13 Galaxy-weighted average $odds$ parameter odds
Figure 3: Result of running our systematics amelioration pipeline on our set of already “processed” J-PLUS-DR3 ADFF maps from the BANNJOS catalogue. See text for details on the non-zero $\epsilon_{i}$ entries for the efectime and texposed templates.
@@
We test the level of residuals remaining in our “corrected” maps from the BANNJOS catalogue by running again our systematics pipeline on them. The result of this test is shown in Fig. 3, and leaves practically no evidence for systematics residuals with the exception of the efectime and texposed templates, which yield opposite entries for the $\epsilon_{i}$ additive parameters in practically every redshift shell. Provided these two templates are extremely similar (and thus highly correlated), it is fair to take the average of their entries, which are very close to zero and carry little to negligible statistical significance. Repeating this same analysis on the “corrected” maps from the XGBoost catalogue gives no significant hint of residuals, for any template nor redshift shell or width.
@@
A particular case is the one associated to the odds template, which, according to the rightmost columns in Fig. 2, presents consistent positive values for multiplicative factor $\beta_{i}$ at the level of $10~{}\%$ for practically all redshift shells. This potential source of systematics was carefully addressed in Martí et al. (2014a), who found odds to be impacting strongly the effective galaxy bias. One must note that the odds parameter is typically higher for brighter galaxies for which photo-$z$ estimates have higher quality. As a result high-odds galaxies reside in high mass halos whose bias is typically higher than a generic galaxy population at the same redshift. While in the case of Martí et al. (2014a) the impact of the odds parameter was corrected at the level of the 2-point statistics, in our case we are attempting its correction at the map level. But in either case the following concern arises: provided that the odds parameter is correlated to a physical property of the galaxies (luminosity or mass), by correcting for it we may be (artificially) modifying the inferred physical bias of our sources. To some extent this in unavoidable provided our basic, underlying assumption about systematics templates is that they are not correlated to the physics of the real galaxy sample, something we expect to (at least partially) fail for the odds parameter. This must be kept in mind, even if the (possible) modulation of the bias seems rather constant throughout redshift bins, and always at the level of $\sim+10~{}\%$.
@@ 5 Results
The joint ADF$+$ARF fits (red curves) are in many cases indistinguishable from the ADF- and ARF-only fits, particularly at low redshifts. Some tension between the 1- and 2-probe fits starts to appear at $z\gtrsim 0.17$, at a higher level for ARF than for ADF. Nevertheless, similar best-fit curves for one- and two-probes do not necessarily involve similar best-fit parameters. Indeed, as can be seen in Fig. 10 for the case where $\sigma_{\rm Err}$ is constant for all redshifts and the BANNJOS selection, the ADF-only and ARF-only contours (blue and green colours, respectively) are clearly incompatible for $\sigma_{\rm Err}$ (the first column of this plot corresponds to this parameter). The tension between the bias contours grows with redshift, surpassing the $\sim 2-\sigma$ level by first time at $z=0.17$, and then again at $z=0.23$ and $z=0.25$. The bias values from ADF at $z<0.17$ are systematically higher than those from ARF, but at the same time, $\sigma_{\rm Err}$ is also much higher. These two parameters are obviously correlated, since higher values of the bias $b_{g}(z)$ combined with higher values of $\sigma_{\rm Err}$ can yield similar amplitudes of the $C_{\ell}$s than those obtained under lower values of these two parameters, provided that each of them have opposite effect on the $C_{\ell}$ amplitudes. Again for the BANNJOS catalogue, Fig 11 shows the projected best-fit values for the parameters under the first parameter configuration (constant $\sigma_{\rm Err}$). The joint ADF$+$ARF fit (in green colour) is driven by the ARF on $\sigma_{\rm Err}$ (since ARF are significantly more sensitive to this parameter), and also for bias parameters below $z<0.19$. Above this redshift the ADF$+$ARF values are dominated by the ADF. The value provided for $\sigma_{\rm Err}$ by the photo-$z$ code LePhare is $\sigma_{\rm Err}=0.010$, which is below (at the $\simeq 9~{}\sigma$ level) the distribution for this parameter from ARF, and also clearly discrepant with the value provided by ADF. Regarding the velocity amplitude ${\cal A}_{v}$, there is a significant mismatch between the ADF and the ARF outputs: ARF see no evidence for radial peculiar velocities (${\cal A}_{v}=0.42\pm 0.57$), ADF point to ${\cal A}_{v}=3.4\pm 1.0$, while the joint ADF$+$ARF fit provides an intermediate solution that just by coincidence peaks around ${\cal A}_{v}\sim 1$, ${\cal A}_{v}=0.90\pm 0.56$.
@@
If we instead look at the results from the second configuration ($\sigma_{\rm Err}$ as a function of redshift), we find again for the BANNJOS catalogue that there is actually agreement for $\sigma_{\rm Err}$ at low redshifts between both probes, although systematically higher for ADF than for ARF, which again translates into higher ADF bias estimates. This is shown in the corner plot Fig. 12, whose results are summarised in Fig. 13. A careful look at Fig. 12 reveals a positive correlation of the $\{\sigma_{\rm Err}(z),b_{g}(z)\}$ parameters for the same shells: this is found in a diagonal whose first panel is located in the first column and 12th row, displaying the correlation between $\sigma_{\rm Err}(z)$ and $b(z)$ for the first redshift bin. This diagonal follows further in the 13th row and second column. In Fig. 13 one can also notice that it is only at the highest redshift shells ($z>0.15$) when $\sigma_{\rm Err}$ estimates for ADF reach the $\sim 0.025$ level, which is higher than any ARF $\sigma_{\rm Err}$ estimate. Thus, on what regards $\sigma_{\rm Err}$ and $b_{g}(z)$, ADF and ARF agree for $z<0.17$, but it must be stressed that ARF provide much tighter constraints on these two parameters, very significantly shrinking their degeneracy area.
@@ 7 Conclusions
The estimates of $\sigma_{\rm Err}(z)$ are compatible from either the ADF or the ARF probe for $z<0.17$ (but almost always higher for ADF than for ARF), a redshift range where we show systematics to have a low impact on the angular distribution of galaxies. These estimates of $\sigma_{\rm Err}(z)$ scatter around $\sigma_{\rm Err}(z)=0.014$ with a typical uncertainty of about 30 %, while the estimate from LePhare is $\langle\sigma_{\rm Err}(z)\rangle=0.010$. ARF are significantly shrinking the degeneracy area between $\sigma_{\rm Err}$ and $b_{g}(z)$, and this translates into more precise estimates of these parameters. On that same redshift range, the galaxy bias estimates from both ADF and ARF seem to grow monotonically with $z$, with a very similar slope, from $b(z)\simeq 0.7$ at $z=0.05$ up to $b(z)\simeq 1.6$ at $z=0.15$. Although compatible, ADF bias estimates again lie systematically above ARF ones, and this is understood given the correlation between $b_{g}(z)$ and $\sigma_{\rm Err}(z)$ for any given redshift shell: higher values of these two parameters impact the amplitude of the angular power spectrum in opposite ways, and it is thus possible to leave the amplitude of the $C_{\ell}$s practically untouched after increasing both parameters conveniently.
@@ acknowledgements
C.H.-M. acknowledges the hospitality of the Centro de Estudios de Física del Cosmos de Aragón (ceFca), where part of this work took place. C.H.-M. also acknowledges the support of the Spanish Ministry of Science and Innovation projects PID2021-126616NB-I00, PID2022-142142NB-I00, the European Union through the grant “UNDARK” of the Widening participation and spreading excellence programme (project number 101159929), and the contribution from the IAC High-Performance Computing support team and hardware facilities. VM thanks CNPq (Brazil) and FAPES (Brazil) for partial financial support. ET acknowledges funding from the HTM (grant TK202), ETAg (grant PRG1006) and the EU Horizon Europe (EXCOSM, grant No. 101159513). AE acknowledges the financial support from the Spanish Ministry of Science and Innovation and the European Union - NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA. Based on observations made with the JAST80 telescope at the Observatorio Astrofísico de Javalambre (OAJ), in Teruel, owned, managed, and operated by the Centro de Estudios de Física del Cosmos de Aragón. We acknowledge the OAJ Data Processing and Archiving Unit Department(DPAD, Cristóbal-Hornillos et al., 2012) for reducing and calibrating the OAJ data used in this work. Funding for the J-PLUS Project has been provided by the Governments of Spain and Aragón through the Fondo de Inversiones de Teruel; the Aragón Government through the Research Groups E96, E103, E16_17R, E16_20R, and E16_23R; the Spanish Ministry of Science, Innovation and Universities (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grants PID2021-124918NB-C41, PID2021-124918NB-C42, PID2021-124918NA-C43, PID2021-124918NB-C44, PGC2018-097585-B-C21 and PGC2018-097585-B-C22; the Spanish Ministry of Economy and Competitiveness (MINECO) under AYA2015-66211-C2-1-P, AYA2015-66211-C2-2, AYA2012-30789, and ICTS-2009-14; and European FEDER funding (FCDD10-4E-867, FCDD13-4E-2685). The Brazilian agencies FINEP, FAPESP, and the National Observatory of Brazil have also contributed to this project. Some of the results in this paper have been derived using the HEALPix (Górski et al., 2005) package.